if you earn an annual interest rate of 10.2 percent how many years will it take to triple your money
Answer #1

FV = PV (1+r)^n

3 = 1 (1+0.102)^n

3 = 1 ( 1.102)^n

1.102^n = 3

Take Log on both sides

Log [ 1.102^n ] = Log 3

Log [ a^b ]

= b * Log [ a ]

n * Log [ 1.102 ] = Log 3

n * 0.0422 = 0.4771

n = 0.4771 / 0.0422

= 11.31

it takes 11.31 Years to triple the value.